What is the pH of the buffer solution in question 2 (Recall the values were: 1.0 L buffer of 0.35 M hydrofluoric acid (HF, Ka = 3.5x10-4) and 0.68 M sodium fluoride (NaF)) after 0.052 moles of HCl are added?

The answer is NOT 2.65
Answer #1

pH of a buffer solution is given by Henderson's equation

pH = pka + logo Salt Acid

Molarity is the number of moles of solute present in one litre of solution.

a- Numberofmolesof solute Volumeof solutionin Liters

Number of moles of solute = M*Volume of solution in L

Given that Ka of HF = 3.5x104 pka = -loska = -log(8.5x104) = 4 - log35 = 4 -0.544 = 3.456 No. of mola of HE in IL of 0.35M HE

When HCl is added to buffer, F​​​​​- ions from NaF interacts with H+ ions forms HF. Therefore number of moles of HF increases and number of moles of NaF decreases.

In the given problem, 0.052 moles of HCl is added.

Therefore number of moles of HF increases by 0.052.

Number of moles of HF after addition of HCl

= 0.35+0.052 = 0.402 moles

Number of moles of NaF decreases by 0.052.

Number of moles of NaF after addition of HCl is

0.68 - 0.052 = 0.628 moles

Since volume of buffer is 1 L

Final concentration of HF = 0.402 M

Final concentration of NaF = 0.628 M

pH = pKa + log([NaF]/[HF])

pH = 3.456+log(0.628/0.402)

pH = 3.456+log(1.562)

pH = 3.456+0.194

pH = 3.65